# Multivariate generalized Pareto distributions along extreme directions

### v3

@@ 5.1 Generic simulation algorithm
Recall the mixture model $\boldsymbol{M}$ in Definition 4.1. The goal of this section is to generate a random variable $\boldsymbol{Y}$ from the associated mgp distribution $H$ in (4.5).
@@
$f_{\boldsymbol{T}^{(k)}}(\boldsymbol{t})=c^{(k)}\,g_{k}(\boldsymbol{t})\,\sum_{j\in J_{k}}n_{j,k}\,q_{j,k}(\boldsymbol{t}),\qquad\boldsymbol{t}\in\mathbb{R}^{\lvert J_{k}\rvert},$ (5.2)
@@
The complexity of Algorithm 1 is driven by the expected number of times one has to simulate from densities $\left({q_{j,k}}\right)_{k\in R,j\in J_{k}}$ in (5.4) and is given by $d/\ell(\boldsymbol{1})$ with $\ell$ in (4.4); see Appendix C for the proof. This is equal to the complexity of the algorithm in Engelke and Hitz (2020, Subsection 5.4) based on extremal functions $\boldsymbol{\mathcal{U}}^{(1)},\ldots,\boldsymbol{\mathcal{U}}^{(d)}$ associated to the mgp random vector $\boldsymbol{Y}$ as defined in Dombry et al. (2016) and characterized in Engelke and Hitz (2020, Lemma 2) by
@@ 5.2.1 Mixture logistic model
We want to simulate the mgp random vector $\boldsymbol{Y}$ associated with the mixture logistic model defined in Section 4.2.1. Let $\boldsymbol{U}$ be the mixture generator in (4.7) with matrix $A$ and generators $\boldsymbol{U}^{(1)},\ldots,\boldsymbol{U}^{(r)}$ in (4.11) with parameters $\alpha_{1},\ldots,\alpha_{r}$. The $\lvert J_{k}\rvert$-variate random vector $\boldsymbol{U}^{(k)}$ has independent Gumbel components and is therefore absolutely continuous with respect to $\lambda_{\lvert J_{k}\rvert}$. Using Algorithm 1, it remains to generate random samples from the densities $q_{j,k}$ ($j\in J_{k}$) defined for $\boldsymbol{t}=\left({t_{i}}\right)_{i\in J_{k}}\in\mathbb{R}^{\lvert J_{k}\rvert}$ by
@@ 5.2.2 Mixture Hüsler–Reiss model
In the same way, we want to draw random samples from the mgp random vector $\boldsymbol{Y}$ associated with the mixture Hüsler–Reiss model in Section 4.2.2. Let $\boldsymbol{U}$ be the mixture generator in (4.7) with matrix $A$ and generators $\boldsymbol{U}^{(1)},\ldots,\boldsymbol{U}^{(r)}$ in (4.14). Let $k\in R$. The $\lvert J_{k}\rvert$-variate random vector $\boldsymbol{U}^{(k)}$ has normal density $f_{\boldsymbol{U}^{(k)}}$. Using Algorithm 1, it remains, for all $j\in J_{k}$, to simulate random samples from the density
$q_{j,k}(\boldsymbol{t})=\frac{\mathrm{e}^{t_{j}}f_{\boldsymbol{U}^{(k)}}\left({\boldsymbol{t}-\left({\ln\left({a_{ik}/m_{k}}\right)}\right)_{i\in J_{k}}}\right)}{a_{jk}/m_{k}},\qquad\boldsymbol{t}=\left({t_{i}}\right)_{i\in J_{k}}\in\mathbb{R}^{\lvert J_{k}\rvert}.$ (5.6)
@@
Let $j\in J_{k}$. Let $\boldsymbol{\zeta}=\left({\zeta_{j}}\right)_{i\in J_{k}}$ have density $q_{j,k}$ in (5.6). With $\tilde{\Sigma}^{(k)}$ as in (4.14), we have
@@ 5.2.3 Simulation example
We use Algorithm 1 with inputs $d=r=3$, triangular coefficient matrix
@@ 6 Conclusion
Additionally, our research opens doors to the application of mgp mixture models in the realm of graphical models for extremes. Whereas the foundations laid by Engelke and Hitz (2020) involve a single extreme direction, the extension offered in Engelke et al. (2022) incorporates general exponent measures, enabling a more comprehensive understanding of extreme events in diverse contexts.
@@ Appendix B Proofs of Section 4
Recall $\mu$, the exponent measure of $G$. For $k\in R$, let $\mu_{J_{k}}^{(k)}$ be the exponent measure of the factor $\boldsymbol{Z}_{J_{k}}^{(k)}$, let $\pi_{J_{k}}:\mathbb{R}^{d}\to\mathbb{R}^{\lvert J_{k}\rvert}$ be the projection $\boldsymbol{x}\mapsto\boldsymbol{x}_{J_{k}}$ and write $\mathbb{F}_{J_{k}}:=\left\{{\boldsymbol{z}\in\mathbb{E}:\;\forall j\notin J_{k},z_{j}=0}\right\}$. For $\boldsymbol{v}\in(0,1)^{m}$ and $A\subseteq[0,\infty)^{m}$, we write $\boldsymbol{v}\,\cdot\,A:=\left\{{\left({v_{1}a_{1},\ldots,v_{m}a_{m}}\right):\ \boldsymbol{a}\in A}\right\}$. Consider the measure
@@
For a non-empty set $J\subseteq D$, recall the set $\mathbb{A}_{J}$ defined in (2.1) and the projection $\pi_{J}:\mathbb{A}_{J}\to\mathbb{R}^{\lvert J\rvert}$ defined via $\boldsymbol{x}\mapsto\boldsymbol{x}_{J}$. Thus, for $J\subseteq D$ and $\boldsymbol{x}\in\mathbb{R}^{\lvert J\rvert}$, the point $\pi^{-1}_{J}(\boldsymbol{x})$ in $\mathbb{A}_{J}$ is defined by setting the components with indices $j\in J$ equal to $x_{j}$ and the components with indices $j\in D\setminus J$ equal to $\infty$- ∞.
@@
Next, let $\boldsymbol{y}\in[0,\infty)^{d}$. Applying (B.2) to the function $g(\boldsymbol{x})=\max\left\{{y_{1}\mathrm{e}^{x_{1}},\ldots,y_{d}\mathrm{e}^{x_{d}}}\right\}$ gives
@@
with $\pi_{J_{k}}:\mathbb{E}_{J_{k}}\to(0,\infty)^{\lvert J_{k}\rvert}$ the natural coordinate projection. Let $\boldsymbol{Z}^{(k)}_{J_{k}}$ be a random vector on $\mathbb{R}^{\lvert J_{k}\rvert}$ with distribution
$\operatorname{\mathsf{P}}\left[{\boldsymbol{Z}^{(k)}_{J_{k}}\leqslant\boldsymbol{x}}\right]=\exp\left\{{-\mu_{J_{k}}^{(k)}\left({[\boldsymbol{0},\boldsymbol{x}]^{c}}\right)}\right\},\qquad\boldsymbol{x}\in[0,\infty]^{\lvert J_{k}\rvert}\setminus\left\{{\boldsymbol{0}}\right\}.$
@@
For $k\in R$, the random vector $\boldsymbol{U}^{(k)}$ in (4.11) satisfies (2.14) and (2.15) for the logistic stdf with parameter $\alpha_{k}$ (Falk, 2019, Proposition 1.2.1). Finally, Proposition 4.7 finishes the proof. ∎
@@
For $k\in R$, the random vector $\boldsymbol{U}^{(k)}$ in (4.11) has independent Gumbel components. Thus $\boldsymbol{U}^{(k)}$ is absolutely continuous with respect to $\lambda_{\lvert J_{k}\rvert}$. Proposition 4.11 yields that the mgp random vector $\boldsymbol{Y}$ associated with the mixture logistic model is absolutely continuous with respect to $\upsilon$ in (3.9) with density $h_{\boldsymbol{Y}}$ given by (4.8). We use Kiriliouk et al. (2018, Subsection 7.2) to calculate the integrals in the latter equation. This yields the density of $\boldsymbol{Y}$ in (4.12). ∎
@@
For $k\in R$, the random vector $\boldsymbol{U}^{(k)}$ in (4.14) satisfies (2.14) and (2.15) for the Hüsler–Reiss stdf with variogram matrix $\Gamma^{(k)}$; see, e.g., Wadsworth and Tawn (2014). Finally, Proposition 4.7 finishes the proof. ∎
@@
For $k\in R$, the random vector $\boldsymbol{U}^{(k)}$ in (4.14) is a $\lvert J_{k}\rvert$-variate normal random vector with positive definite covariance matrix. Thus $\boldsymbol{U}^{(k)}$ is absolutely continuous w.r.t. $\lambda_{\lvert J_{k}\rvert}$. Proposition 4.11 yields that the mgp random vector $\boldsymbol{Y}$ associated with the mixture Hüsler–Reiss model is absolutely continuous with respect to $\upsilon$ with density $h_{\boldsymbol{Y}}$ given by (4.8). We use Kiriliouk et al. (2018, Section 7.2) to calculate the integrals in the latter equation. This yields the density of $\boldsymbol{Y}$ given by (4.15). ∎
@@ Appendix C Proofs of Section 5
where $\boldsymbol{T}^{(k)}$ is a random vector on $\mathbb{R}^{\lvert J_{k}\rvert}$ which satisfies (5.1).
@@
where the last equality is proved in Engelke and Hitz (2020, Lemma 2). ∎

### v4

@@ 5.1 Generic simulation algorithm
Recall the mixture model $\boldsymbol{M}$ in Definition 4.1. The goal of this section is to generate a random variable $\boldsymbol{Y}$ from the associated mgp distribution $H$ in (4.5) and with density in Proposition 4.11.
@@
$f_{\boldsymbol{T}^{(k)}}(\boldsymbol{t})=c^{(k)}\,g_{k}(\boldsymbol{t})\,\sum_{j\in J_{k}}n_{j,k}\,q_{j,k}(\boldsymbol{t}),\qquad\boldsymbol{t}\in\mathbb{R}^{J_{k}},$ (5.2)
@@
The complexity of Algorithm 1 is driven by the expected number of times one has to simulate from densities $\left({q_{j,k}}\right)_{k\in R,j\in J_{k}}$ in (5.4) and is given by $d/\ell(\boldsymbol{1})$ with $\ell$ in (4.4); see Appendix C for the proof. This is equal to the complexity of the algorithm in Engelke and Hitz [16, Subsection 5.4] based on extremal functions $\boldsymbol{\mathcal{U}}^{(1)},\ldots,\boldsymbol{\mathcal{U}}^{(d)}$ as defined in Dombry et al. [12] and characterized in Engelke and Hitz [16, Lemma 2] in terms of the mgp random vector $\boldsymbol{Y}$ by
@@ 5.2.1 Mixture logistic model
We want to simulate the mgp random vector $\boldsymbol{Y}$ associated with the mixture logistic model defined in Section 4.2.1. Let $\boldsymbol{U}$ be the mixture generator in (4.7) with matrix $A$ and generators $\boldsymbol{U}^{(1)},\ldots,\boldsymbol{U}^{(r)}$ in (4.11) with parameters $\alpha_{1},\ldots,\alpha_{r}$. The $\lvert J_{k}\rvert$-variate random vector $\boldsymbol{U}^{(k)}$ has independent Gumbel components and is therefore absolutely continuous with respect to $\lambda_{\lvert J_{k}\rvert}$. Using Algorithm 1, it remains to generate random samples from the densities $q_{j,k}$ ($j\in J_{k}$) defined for $\boldsymbol{t}=\left({t_{i}}\right)_{i\in J_{k}}\in\mathbb{R}^{J_{k}}$ by
@@ 5.2.2 Mixture Hüsler–Reiss model
In the same way, we want to draw random samples from the mgp random vector $\boldsymbol{Y}$ associated with the mixture Hüsler–Reiss model in Section 4.2.2. Let $\boldsymbol{U}$ be the mixture generator in (4.7) with matrix $A$ and generators $\boldsymbol{U}^{(1)},\ldots,\boldsymbol{U}^{(r)}$ in (4.14). Let $k\in R$. The $\lvert J_{k}\rvert$-variate random vector $\boldsymbol{U}^{(k)}$ has normal density $f_{\boldsymbol{U}^{(k)}}$. Using Algorithm 1, it remains to simulate from the densities $q_{j,k}$ ($j\in J_{k}$) in Equation (5.5), with the difference that the generators $\boldsymbol{U}^{(1)},\ldots,\boldsymbol{U}^{(r)}$ are defined via (4.14) instead of (4.11).
@@
Let $j\in J_{k}$. Let $\boldsymbol{\zeta}=\left({\zeta_{j}}\right)_{i\in J_{k}}$ have density $q_{j,k}$ in (5.5) with generators $\boldsymbol{U}^{(1)},\ldots,\boldsymbol{U}^{(r)}$ as in (4.14). With $\tilde{\Sigma}^{(k)}$ as in (4.14), we have
@@ 5.2.3 Simulation example
The results in this section can be reproduced using the code available on https://github.com/AnasMourahib/MGPD-simulation. We use Algorithm 1 with inputs $d=r=3$, triangular coefficient matrix
@@ 6 Conclusion
Additionally, our research opens doors to the application of mgp mixture models in the realm of graphical models for extremes. Whereas the foundations laid by Engelke and Hitz [16] involve a single extreme direction, the extension offered in Engelke et al. [18] incorporates general exponent measures, enabling a more comprehensive understanding of extreme events in diverse contexts.
@@ Appendix B Proofs of Section 4
Recall $\mu$, the exponent measure of $G$. For $k\in R$, let $\mu_{J_{k}}^{(k)}$ be the exponent measure of the factor $\boldsymbol{Z}_{J_{k}}^{(k)}$, let $\pi_{J_{k}}:\mathbb{R}^{d}\to\mathbb{R}^{J_{k}}$ be the projection $\boldsymbol{x}\mapsto\boldsymbol{x}_{J_{k}}$ and write $\mathbb{F}_{J_{k}}:=\left\{{\boldsymbol{z}\in\mathbb{E}:\;\forall j\notin J_{k},z_{j}=0}\right\}$. For $\boldsymbol{v}\in(0,1)^{m}$ and $A\subseteq[0,\infty)^{m}$, we write $\boldsymbol{v}\,\cdot\,A:=\left\{{\left({v_{1}a_{1},\ldots,v_{m}a_{m}}\right):\ \boldsymbol{a}\in A}\right\}$. Consider the measure
@@
For a non-empty set $J\subseteq D$, recall the set $\mathbb{A}_{J}$ defined in (2.1) and the projection $\pi_{J}:\mathbb{A}_{J}\to\mathbb{R}^{J}$ defined via $\boldsymbol{x}\mapsto\boldsymbol{x}_{J}$. Thus, for $J\subseteq D$ and $\boldsymbol{x}\in\mathbb{R}^{J}$, the point $\pi^{-1}_{J}(\boldsymbol{x})$ in $\mathbb{A}_{J}$ is defined by setting the components with indices $j\in J$ equal to $x_{j}$ and the components with indices $j\in D\setminus J$ equal to $\infty$- ∞.
@@
Next, let $\boldsymbol{y}\in[0,\infty)^{d}$. Applying (B.2) to the function $g(\boldsymbol{x})=\max\left({y_{1}\mathrm{e}^{x_{1}},\ldots,y_{d}\mathrm{e}^{x_{d}}}\right)$ gives
@@
with $\pi_{J_{k}}:\mathbb{E}_{J_{k}}\to(0,\infty)^{J_{k}}$ the natural coordinate projection. Let $\boldsymbol{Z}^{(k)}_{J_{k}}$ be a random vector on $\mathbb{R}^{J_{k}}$ with distribution
$\operatorname{\mathsf{P}}\left[{\boldsymbol{Z}^{(k)}_{J_{k}}\leqslant\boldsymbol{x}}\right]=\exp\left\{{-\mu_{J_{k}}^{(k)}\left({[\boldsymbol{0},\boldsymbol{x}]^{c}}\right)}\right\},\qquad\boldsymbol{x}\in[0,\infty]^{J_{k}}\setminus\left\{{\boldsymbol{0}}\right\}.$
@@
For $k\in R$, the random vector $\boldsymbol{U}^{(k)}$ in (4.11) satisfies (2.14) and (2.15) for the logistic stdf with parameter $\alpha_{k}$ [19, Proposition 1.2.1]. Finally, Proposition 4.7 finishes the proof. ∎
@@
For $k\in R$, the random vector $\boldsymbol{U}^{(k)}$ in (4.11) has independent Gumbel components. Thus $\boldsymbol{U}^{(k)}$ is absolutely continuous with respect to $\lambda_{\lvert J_{k}\rvert}$. Proposition 4.11 yields that the mgp random vector $\boldsymbol{Y}$ associated with the mixture logistic model is absolutely continuous with respect to $\upsilon$ in (3.10) with density $h_{\boldsymbol{Y}}$ given by (4.8). We use Kiriliouk et al. [29, Subsection 7.2] to calculate the integrals in the latter equation. This yields the density of $\boldsymbol{Y}$ in (4.12). ∎
@@
For $k\in R$, the random vector $\boldsymbol{U}^{(k)}$ in (4.14) satisfies (2.14) and (2.15) for the Hüsler–Reiss stdf with variogram matrix $\Gamma^{(k)}$; see, e.g., Wadsworth and Tawn [42]. Finally, Proposition 4.7 finishes the proof. ∎
@@
For $k\in R$, the random vector $\boldsymbol{U}^{(k)}$ in (4.14) is a $\lvert J_{k}\rvert$-variate normal random vector with positive definite covariance matrix. Thus $\boldsymbol{U}^{(k)}$ is absolutely continuous w.r.t. $\lambda_{\lvert J_{k}\rvert}$. Proposition 4.11 yields that the mgp random vector $\boldsymbol{Y}$ associated with the mixture Hüsler–Reiss model is absolutely continuous with respect to $\upsilon$ with density $h_{\boldsymbol{Y}}$ given by (4.8). We use Kiriliouk et al. [29, Section 7.2] to calculate the integrals in the latter equation. This yields the density of $\boldsymbol{Y}$ given by (4.15). ∎
@@ Appendix C Proofs of Section 5
where $\boldsymbol{T}^{(k)}$ is a random vector on $\mathbb{R}^{J_{k}}$ which satisfies (5.1).
@@
where the last equality is proved in Engelke and Hitz [16, Lemma 2]. ∎